OpenStudy (anonymous):

How do you rewrite this expression using positive exponents only? 3s^1/3 * s^-7/3

OpenStudy (anonymous):

3/x^2

OpenStudy (anonymous):

\[3s ^{1/3}\times s ^{-7/3}\]

OpenStudy (anonymous):

\[3s ^{1/3}\times s ^{-7/3} = 3 \times s ^{(1/3) + (-7/3)}\] \[3 \times s ^{-6/3}\] \[3 \times s ^{-2}\] since \[s ^{-2} = 1/s^2\] so we have \[3/s^2\]